What is the mole fraction of phosphoric acid (H3PO4) in a 25.4% (by mass) aqueous solution of the acid? Enter your answer with four decimal places and no units.
Answer #1

25.4% by mass means

25.4 g of H3PO4 in 100g of aqueous solution.

Thus the mass of water in solution = 100- 25.4 =74.6 g

moles of H3PO4 = mass / molar mass

= 25.4g / 98g/mol =0.2591 mol

moles of water in solution = mass of water / molar mass

= 74.6g/ 18 g/mol = 4.14 mol

thus

mole fraction of phosphoric acid = moles of H3PO4 / moles of H3PO4 + moles of H2O

= 0.2591/ ( 0.2591+4.14)

= 0.0588